A right-circural cone is inscribed in a sphere having a fixed radius of 10 in. Express the volume of the cone as a function of its radius.
1
Expert's answer
2020-10-05T16:54:17-0400

volume of the cone = 1/3 "\pi"r2H

r=radius of the cone

H=height of the cone

R= radius of the sphere=10

H=R+(R2-r2)1/2 or H=H=R-(R2-r2)1/2

the function of the volume from its base now =

1/3 "\pi"r2H = 1/3 "\pi"r2*(R+(R2-r2)1/2)

=1/3 "\pi"r2*(10+(100-r2)1/2)

or

1/3 "\pi"r2H = 1/3 "\pi"r2*(R-(R2-r2)1/2)

=1/3 "\pi"r2*(10-(100-r2)1/2)